OpenStudy (anonymous):

Solve for x: |2x + 6| − 4 = 20 x = 9 and x = 11 x = −9 and x = 15 x = 9 and x = −15 No solution

OpenStudy (anonymous):

|2x + 6| - 4 = 20 2x + 6 = 24 & 2x + 6 = -24 2x = 18 & 2x = -30 x = 9 & x = -15 As you can see, you get whatever's in the | x | by itself on one side and set it equal to both the positive and negative solutions. Solve for c.

OpenStudy (anonymous):

solve for x*